Analyzling Manufacturing Cost Accounts Fire Rock Company manufactures designer padle boards in a wide variety of sizes and styles. The folloning Incomplete ledger e boards in a wide variety of sizes and styles. The following incomplete ledger accounts refer to transactions that are summarized for June Materials une I une 30 Balance 0 Requisitions Purchases Work in Process June 1 une 30 Materials June 30 Direct labor Balance (e)bune 30 (C Completed jobs CF) く)re 30 Factory overhead applied (E Finished Goods une i ODune 3 0 Cost of goods sold June 30 Completed jobs Wages Payable une 3 0 Wages incurred 122,500 Factory Overhead une 1 une 30 lune 30 Indirect materi une 30 oner overhead 20,800une 30 Factory overhead applied Indirect labor 18,900 75,900
Print Item In addition, the following information is available: a. Materials and direct labor were applied to six jobs in July: Job No. Style Quantity Direct Materials Direct Labor 201 202 203 204 205 206 $15,000 T100 T200 T400 S200 T300 S100 Total 210 370 170 310 170 120 1,350 $24,780 36,380 13,460 45,650 22,460 8,640 24,000 7,000 32,000 16,000 4,000 $151,370 $98,000 b. Factory overhead is applied to each job at a rate of 140% of direct labor cost. c. The June 1 Work in Process balance consisted of two jobs, as follows: Job No. Style 201 202 Total Work in Process, June 1 $6,000 14,800 T100 T200 $20,800 d. Customer jobs completed and units sold in July were as follows: Job No. Style Completed in July Units Sold in July 201 T100 202 T200 203 T400 204 S200 205 T300 206 S100 168 296 260 141 Check My Work


eBook Calculator Print item Work inMaterials Labor Total Cost ob No. Quantity Unit Cost Units Sold Cost of Goods Sold Overhead Process No. 201 No. 202 No. 203 No. 204 No. 205 No. 206 Total 6,000 $24,780$15,000 14,800 36,380 24,000 13,4607,000 45,65032,000 22,460 16,000 8,640 4,000 $20,800 151,370 98,000 a. Materials Requistions b. Work in Process Beginning Balance c. Direct Materials d. Direct Labor e. Factory overhead applied 1. Completed jobs 9 Cost of goods sold h. Indirect labor 2. Determine the June 30 balances for each of the inventory accounts and factory overhead. Use the mius sign to indicate any credt balances Materials: Work in Process: Finished Goods Factory Overhead:
Answer #1

Calculation of required fields in ledger:

Overhead Rate 140% of direct labor cost FactoryTotal Cost of Cost of Goods Sold Work in Job No Quantity Process Direct Materials Direct Labor OverheadCompleted Jobs Unit cost Units sold 201 202 203 210 $6,000 370 $14,800 170 310 170 4,780 $15,000 $21,000 33,600 $7,000$9,800 45,65032,000 $44,800 $22,460$16,000 $22,400 S4,000 $5,600 $151,370 $98,000 $137,200 $2 36,380 $24,000S3 13,460 $66,780 $108,780 318 $53,424 87,024 168 296 $395 102,700 S50,478 $60,860 141 $8,640 Total $20,800 $358,870 $293,626 12 Materials $170,270 1-Jun Balance 30-Jun Purchases 35,400 30un Requisitions 15 16 142,300 1-Jun Balance 30-Jun Materials 30-Jun Direct Labor 30-Jun Factory overhead applied Work in process $20,80030-Jun Completed jobs $151,370 $98,000 137,200 358,870 Finished Goods 1-Jun Balance 30-Jun Completed Jobs $0 30-Jun Cost of Goods Sold $293,626 27 $358,870 Wages Payable 30Jun Wages Incurred $122,500 31 Factory Overhead 1-Jun Balance 30-Jun Indirect Labor 30-Jun Indirect Materials 30-Jun Other Overhead 20,80030-Jun Factory Overhead applied $24,500 $18,900 $75,900 $137,200 35 36 37 Hence Materials Requisitions $170,270 Work in Process Beginning Bal $20,800 $151,370 $98,000 $137,200 $358,870 $293,626 $24,500 Direct Materials Direct Labor Overhead Applied Completed Jobs Cost of Indirect Labor

Formula sheet

A B C D E F G H I J K L M N
2
3 Overhead Rate 1.4 of direct labor cost
4 Job No Quantity Work in Process Direct Materials Direct Labor Factory Overhead Total Cost of Completed Jobs Unit cost Units sold Cost of Goods Sold
5 201 210 6000 24780 15000 =H5*$E$3 =F5+G5+H5+I5 =J5/E5 168 =K5*L5
6 202 370 14800 36380 24000 =H6*$E$3 =F6+G6+H6+I6 =J6/E6 296 =K6*L6
7 203 170 13460 7000 =H7*$E$3 =K7*L7
8 204 310 45650 32000 =H8*$E$3 =F8+G8+H8+I8 =J8/E8 260 =K8*L8
9 205 170 22460 16000 =H9*$E$3 =F9+G9+H9+I9 =J9/E9 141 =K9*L9
10 206 120 8640 4000 =H10*$E$3 =K10*L10
11 Total =SUM(F5:F10) =SUM(G5:G10) =SUM(H5:H10) =SUM(I5:I10) =SUM(J5:J10) =SUM(M5:M10)
12
13
14 Materials
15 43617 Balance 35400 43646 Requisitions =E21+E36
16 43646 Purchases 142300
17
18
19 Work in process
20 43617 Balance =F11 43646 Completed jobs =J11
21 43646 Materials =G11
22 43646 Direct Labor =H11
23 43646 Factory overhead applied =I11
24
25 Finished Goods
26 43617 Balance =0 43646 Cost of Goods Sold =M11
27 43646 Completed Jobs =H20
28
29 Wages Payable
30 43646 Wages Incurred 122500
31
32
33 Factory Overhead
34 43617 Balance 20800 43646 Factory Overhead applied =E23
35 43646 Indirect Labor =H30-H11
36 43646 Indirect Materials 18900
37 43646 Other Overhead 75900
38
39 Hence
40 a) Materials Requisitions =H15
41 b) Work in Process Beginning Bal =E20
42 c) Direct Materials =E21
43 d) Direct Labor =E22
44 e) Factory Overhead Applied =E23
45 f) Completed Jobs =H20
46 g) Cost of goods sold =H26
47 h) Indirect Labor =E35
48